OpenStudy (anonymous):

solve for x 4(x+16)=-4(x-24)

OpenStudy (anonymous):

@XxdeadshotxX 4 (x + 16) = -4(x - 24) 4x + 64 = -4x + 96 4x + 64 + 4x = -4x + 96 + 4x 8x + 64 = 96 8x + 64 - 64 = 96 - 64 8x = 32 8x/8 = 32/8 x = 4

jagr2713 (jagr2713):

@cutelilcigarette please explain next time

OpenStudy (anonymous):

Like say you're distributing and stuff

OpenStudy (anonymous):

I double checked it and it's all good.

OpenStudy (anonymous):

@XxdeadshotxX 4 (x + 16) = -4(x - 24) simplify both sides of the equation 4x + 64 = -4x + 96 add 4x to both sides 4x + 64 + 4x = -4x + 96 + 4x 8x + 64 = 96 subtract 64 from both sides 8x + 64 - 64 = 96 - 64 8x = 32 divide both sides by 8 8x/8 = 32/8 and you get x = 4 better? lol

jagr2713 (jagr2713):

@cutelilcigarette Thanks!

OpenStudy (anonymous):

welcome! @jagr2713